The highway fuel economy of a 2016 Lexus RX 350 FWD 6-cylinder 3.5-L automatic 5-speed using premium fuel is a normally distributed random variable with a mean of μ = 24.50 mpg and a standard deviation of σ = 3.50 mpg.

(a) What is the standard error of X¯¯¯ , the mean from a random sample of 9 fill-ups by one driver?

(b) Within what interval would you expect the sample mean to fall, with 95 percent probability? (Round your answers to 4 decimal places.)
  
The interval is from ___ to ___
Answer #1

A = 24:50, v = 3 5 6 @ standard erro ,se - ne a = 3:50 = 1.1667). 3.50 ca 95% = 0.95 → a=140.95 = 0.05 2/2 = 0.025 + Zo = 20.